# Linguistic Collapse: Neural Collapse in (Large) Language Models

### v1

@@ 3.6 Alignment Between Classifiers and Class Mean Embeddings – $\mathcal{(U)NC}_{3}$
We define uniform duality ($\mathcal{UNC}_{3}$) to be the minimization of the CoV of these similarities.
3.7 Agreement of the Classifiers – $\mathcal{NC}_{4}$
@@ 4 Results
4.1 Within-Class Variability Collapse – $\mathcal{NC}_{1}$
Scaling our models dramatically reduces normalized variance, which is further aided by more training epochs and stronger weight decay (Appendix Figs. 5, 6). These noise reductions tightly associate with generalization (Fig. 1, left, “$\mathcal{NC}_{1}$”). The relationship is strongest at larger model scale.
4.2 Geometric Structures – $\mathcal{(G)NC}_{2}$
Figure 2: Validation loss shows a negligible correlation with equiangularity ($\mathcal{NC}_{2}$, left) but a strong one with hyperspherical uniformity ($\mathcal{GNC}_{2}$, right). In other words, $\mathcal{GNC}_{2}$ develops with scale and correlates well with generalization, much better than $\mathcal{NC}_{2}$.
@@ 4.2 Geometric Structures – $\mathcal{(G)NC}_{2}$
Mean norms grow with model width and training (Appendix Fig. 7), and subtly with depth (Appendix Fig. 8). Despite this growth, the variation of norms consistently decreases with scale (Appendix Figs. 9, 10). Both trends correlate with improved generalization (Appendix Fig. 25).
@@
Scaling model dimensions reduces average interference (Appendix Figs. 11, 12) down to an empirical plateau of approximately $10^{-3}$, which is more apparent in less-trained models. However, the variation of interference rises and persists when scaling (Appendix Figs. 13, 14), suggesting that interference becomes more concentrated between some pairs of classes. These results could be due to various factors, including but not limited to unfriendly conditions of language modeling (§1.2) or the impossibility of a simplex ETF (§3.5).
@@
Logarithmic distances drop more gradually and consistently with scale (Appendix Figs. 15, 16), implying this quantity is more robust or may not face the same geometric barriers seen in conventional interference (Appendix Figs. 13, 14). Variation of these logarithmic distances is also consistently reduced with scale (Appendix Fig. 17, 18).
@@
4.3 Classifier (Mis)alignment and Duality – $\mathcal{(U)NC}_{3}$
@@
Figure 3: Validation loss shows a negligible relationship with self-duality ($\mathcal{NC}_{3}$, left) and some correlation with uniform duality ($\mathcal{UNC}_{3}$, right). In other words, $\mathcal{UNC}_{3}$ develops with scale and correlates with generalization well and much better than $\mathcal{NC}_{3}$.
4.4 Classifier Agreement – $\mathcal{NC}_{4}$
The linear and NCC classifiers agree on far more examples than random chance. Scaling encourages agreement (Appendix Figs. 23, 24). Increasing width for certain depths happens to plateau or even regress the agreement rate, but this limitation is overcome with further training. And finally, agreement is a strong indicator of generalization (Fig. 1, right, $\mathcal{NC}_{4}$).
@@ 4.7 The Effect of Weight Regularization
Our models trained with either weight decay factor exhibited very similar patterns in the emergence of $\mathcal{NC}$ (or lack thereof), but the more aggressive factor $\beta=0.1$ typically used to train LLMs [77] resulted in stronger development of $\mathcal{NC}$ properties (Appendices D, E, F, G, H, I, J, K, L, M).
@@
$\mathcal{NC}$1 Within-Class Variability Collapse $0.192011$ $\mathbf{0.0485}$
$\mathcal{NC}$2 Equinormness $0.026174$ $0.4870$
$\mathcal{NC}$2 Equiangularity $0.218574$ $\mathbf{0.0317}$
$\mathcal{GNC}$2 Hyperspherical Uniformity $0.487935$ $\mathbf{0.0002}$
$\mathcal{NC}$3 Self-Duality $0.322210$ $\mathbf{0.0063}$
$\mathcal{UNC}$3 Uniform Duality $0.000036$ $0.9784$
$\mathcal{NC}$4 Classifier Agreement $0.490278$ $\mathbf{0.0001}$
@@ 5 Limitations
While to the best of our knowledge, no previous work has studied realistic stochastic token prediction, it is possible that the quantities that we measure are not perfectly suited for $\mathcal{NC}$ in language modeling.
Our work is on causal language modeling in its most basic form. We did not conduct experiments into encoding, multi-modal, or instruction-tuned models.
The models that we used in our permutation test (§4.5, Table 1) are only of single small architecture. Therefore, our results on scale-independent $\mathcal{NC}$ do not necessarily translate to larger models.
@@ 6 Discussion
Deep neural collapse ($\mathcal{DNC}$)
A large body of work has established that properties resembling $\mathcal{NC}$ evolve as a function of model depth [33, 52, 58, 59, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101]. Our results mostly concur, but encourage further exploration of $\mathcal{DNC}$ in language modeling.
@@
Given the evidence for the development of $\mathcal{NC}$ and associated generalization under various loss functions [16, 48, 59, 97, 102, 103, 104] in other domains, NLP researchers may also benefit from analyzing, adapting or training towards $\mathcal{NC}$ with respect to alternative objectives.
Compression
CLMs appear to reduce noise, demonstrating their ability to encapsulate different textual contexts into single embeddings. This subscribes to a theory that LLMs are essentially a compression of internet content by aleatoric modeling [105]. However, owing to contextual ambiguity, we hypothesize that CE loss and noise (Equation 4) cannot be fully eliminated by current practices.
@@
Our formulae (Section 3) and results (Section 4) expose the pair-wise token class interactions in noise, interference, classifier duality, and classifier agreement in the top-level features. Similarly to works in other domains [18, 54, 70, 106], these $\mathcal{NC}$ metrics can serve as a form of low-level interpretability to aid understanding certain behaviors of (L)LMs.
@@
Foundation models are ubiquitous for their comprehensive capabilities and adaptability. As previous work discussed class imbalance [32, 57, 60, 61], our work may extend these strategies to measure and perhaps promote fairness in foundation LLMs, some of which are designed to be multilingual or multicultural. While $\mathcal{NC}$ itself would not lead to unfairness, its potential interpretability may, in theory, enable an adversarial agent to measure (un)fairness as they manipulate a LLM.
@@ 7 Conclusion
In this paper, we adapt the neural collapse ($\mathcal{NC}$) framework to the language modeling domain, where models are undertrained and next-tokens are variably drawn from numerous and imbalanced token classes. We leverage the canonical and more recent metrics to demonstrate that $\mathcal{NC}$ emerges as we scale the size and training of hundreds of causal language models. Our results show a correlation between $\mathcal{NC}$ and generalization, a relationship that persists even when the model scale is fixed.
With these novel insights, we hope to inspire new efforts to study $\mathcal{NC}$ (and related behaviors) in language modeling and capitalize on their previously established benefits to improve and better understand the increasingly complex (large) language models and their training processes.
Appendix A Architectural Details
@@
Appendix B Dataset
TinyStories is a synthetic dataset of short children’s stories generated by GPT-3.5 and GPT-4 [2].
  • The dataset is released with the CDLA-Sharing-1.0 licence.
  • Our models were trained and evaluated using the first version of TinyStories [2].
  • Their experimental setup [2] called for the GPT-2 [66] tokenizer, of which only a subset vocabulary $\mathcal{V}$ of size $29,233$ appears in TinyStories.
[…]
  • Dataloading and preprocessing (including chunking) each used two workers.
Figure 4: The $500$ most frequent classes from TinyStories [2] exhibit significant sample imbalance.
@@ Appendix C Optimization
Appendix D Within-Class Variability Collapse with Scale – $\mathcal{NC}$1
Figure 5: Average (logarithmic) class-distance normalized variance (CDNV) is reduced ($\mathcal{NC}$1) when scaling width ($d$) and across training for 1 (left) through 10 (right) epochs with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Figure 6: Average (logarithmic) class-distance normalized variance (CDNV) is reduced ($\mathcal{NC}$1) when scaling depth ($L$) and across training for 1 (left) through 10 (right) epochs with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Appendix E Mean Norms Growth with Scale – (Related to $\mathcal{NC}$2)
@@
Appendix F Equinormness with Scale – $\mathcal{NC}$2
Figure 9: Variation in embedding norms decreases ($\mathcal{NC}$2) when scaling width ($d$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). Note that the degree of equinormness eventually plateaus for sufficiently deep and trained models.
Figure 10: Variation in embedding norms decreases ($\mathcal{NC}$2,) when scaling depth ($L$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Appendix G Interference with Scale – (Related to $\mathcal{NC}$2)

### v2

@@ 3.6 Alignment Between Classifiers and Class Mean Embeddings – $\mathcal{(U)NC}3$
For intuition analogous to that for equinormness and equiangularity (§3.5), we also measure uniform duality ($\mathcal{UNC}_{3}$) as the minimization of the CoV of these similarities (Appendices N, O).
3.7 Agreement of the Classifiers – $\mathcal{NC}4$
@@ 4 Experimental Results
4.1 Within-Class Variability Collapse – $\mathcal{NC}1$
Scaling our models dramatically reduces normalized variance, which is further aided by more training epochs and stronger weight decay (Appendix Figs. 6, 7). These noise reductions tightly associate with generalization (Fig. 1, left, “$\mathcal{NC}_{1}$”). The relationship is most apparent at larger model scale.
4.2 Geometric Structures – $\mathcal{(G)NC}2$
Figure 2: Validation loss shows negligible correlation with equinormness ($\mathcal{NC}_{2}$, left), some relationship with equiangularity ($\mathcal{NC}_{2}$, centre), and a stronger one with hyperspherical uniformity ($\mathcal{GNC}_{2}$, right). So, $\mathcal{GNC}_{2}$ develops with scale and correlates well with generalization, better than $\mathcal{NC}_{2}$.
@@ 4.2 Geometric Structures – $\mathcal{(G)NC}2$
Class mean embedding norms grow with model width and training (Appendix Fig. 8), and subtly with depth (Appendix Fig. 9). Despite this growth, the variation of norms consistently decreases (improving equinormness) with scale (Appendix Figs. 10, 11). Both trends correlate with improved generalization (Appendix Fig. 20).
@@
Scaling model dimensions reduces average interference (Appendix Figs. 12, 13) down to an empirical plateau of approximately $10^{-3}$, which is more apparent in less-trained models. However, the variation of interference rises and persists when scaling (Appendix Figs. 14, 15), suggesting that interference becomes more concentrated between some pairs of classes. These results could be due to various factors, including but not limited to unfriendly conditions of language modeling (§1.2) or the impossibility of a simplex ETF (§3.5).
@@
Logarithmic distances drop more gradually and consistently with scale (Appendix Figs. 16, 17), implying this quantity is more robust or may not face the same geometric barriers seen in conventional interference (Appendix Figs. 14, 15). Variation of these logarithmic distances is also consistently reduced with scale (Appendix Fig. 18, 19).
@@
4.3 Classifier (Mis)alignment and Duality – $\mathcal{(U)NC}3$
@@
Figure 3: Validation loss shows a negligible relationship with self-duality ($\mathcal{NC}_{3}$, left) and some correlation with uniform duality ($\mathcal{UNC}_{3}$, right). In other words, $\mathcal{UNC}_{3}$ develops with scale and correlates with generalization much better than $\mathcal{NC}_{3}$.
4.4 Classifier Agreement – $\mathcal{NC}4$
The linear and NCC classifiers agree on far more examples than random chance. Scaling encourages agreement (Appendix Figs. 29, 30). Increasing width for certain depths happens to plateau or even regress the agreement rate, but this limitation is overcome with further training. And finally, agreement is a strong indicator of generalization (Fig. 1, right, $\mathcal{NC}_{4}$).
5 Analysis
@@ 5.3 The Effect of Weight Regularization
Our models trained with either weight decay factor exhibited very similar patterns in the emergence of $\mathcal{NC}$ (or lack thereof), but the more aggressive factor $\beta=0.1$ resulted in stronger development of $\mathcal{NC}$ properties than with $\beta=0.0005$ (Appendices E, F, G, H, I, J, K, M, N, P). These findings empirically affirm $\beta=0.1$ weight decay as sensible for LLM pre-training [81], and concur with [56] on the pivotal role that appropriate regularization plays in the emergence of $\mathcal{NC}$.
@@
$\mathcal{NC}_{1}$ Within-Class Variability Collapse $0.192011$ $\mathbf{0.0485}$
$\mathcal{NC}_{2}$ Equinormness $0.026174$ $0.4870$
$\mathcal{NC}_{2}$ Equiangularity $0.218574$ $\mathbf{0.0317}$
$\mathcal{GNC}_{2}$ Hyperspherical Uniformity $0.487935$ $\mathbf{0.0002}$
$\mathcal{NC}_{3}$ Self-Duality $0.322210$ $\mathbf{0.0063}$
$\mathcal{UNC}_{3}$ Uniform Duality $0.000036$ $0.9784$
$\mathcal{NC}_{4}$ Classifier Agreement $0.490278$ $\mathbf{0.0001}$
@@ 6 Limitations
Neural collapse
While to the best of our knowledge, no previous work has studied realistic stochastic token prediction, it is possible that the quantities that we measure are not perfectly suited for $\mathcal{NC}$ in language modeling. As we described in §1.1, the $\mathcal{NC}$ framework does not translate neatly to the language modeling space due to many adverse conditions, so full convergence to $\mathcal{NC}$ in the TPT was highly improbable. This paper leaves much room for future work to better adapt $\mathcal{NC}$ for next-token prediction, which we discuss further in Section 7.
Language modeling
Our work focused on autoregressive pre-training in its most basic form. We did not conduct experiments into encoder, multi-modal, or instruction-tuned models. Post-training techniques such as supervised fine-tuning, reinforcement learning with human feedback [91] or direct preference optimization [92] are also out-of-scope. This paper uses validation CE loss as the sole indicator of performance, leaving out any downstream task evaluations.
Confounder of model scale
The models that we used in our permutation test (§5.1, Table 1) are only of a single small architecture trained for one epoch with relatively weak weight regularization ($\beta=0.0005$). Therefore, our experimental results on scale-independent links between $\mathcal{NC}$ and generalization may not necessarily translate to larger models. Further investigation on (foundation) LLMs orders of magnitude larger than our CLMs trained with modern NLP methods would provide more robust insight into any direct correlations.
@@ 7 Discussion
Layer/depth-wise neural collapse
Past works have established that properties resembling $\mathcal{NC}$ evolve as a function of model depth [4, 36, 53, 59, 60, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105]. Layer-wise $\mathcal{NC}$ – sometimes dubbed deep neural collapse ($\mathcal{DNC}$) [53, 59] – and related phenomena at intermediate layers remain an interesting subtopic. We leave their study and induction in CLMs (like [106]) as future work.
@@
Given the evidence for the development of $\mathcal{NC}$ and associated generalization under various loss functions [19, 49, 60, 101, 107, 108, 109] in other domains, NLP researchers may still benefit from analyzing, adapting or training towards $\mathcal{NC}$. As alluded to earlier, the simplex ETF and even the CE loss may not be truly optimal for this problem setting, so we anticipate future works to both construct more amenable geometries with better-suited objectives and then capitalize on their benefits downstream. As discussed in §2, there is an abundance of literature in $\mathcal{NC}$, some of which could potentially adapt $\mathcal{NC}$ to be useful for NLP; we hope to inspire such investigations.
LLM Evaluations
Researchers in NLP and multimodal settings are ultimately interested in measuring model performance on practical tasks. Notable benchmarks include GLUE [77], MMLU [110], and BIG-bench [111]. However, several contemporaries [112, 113, 114] have demonstrated that models’ downstream capabilities are roughly correlated with their abilities to effectively compress their pre-training data. Based on their findings, our application of the $\mathcal{NC}$ framework to the pre-training stage of CLMs against validation CE loss should be an appropriate first step in this intersection. Looking forward, we anticipate exciting analysis for language modeling tasks or benchmarks, especially on creativity and retrieval for natural language understanding and generation.
Conversely, some form of $\mathcal{NC}$ could even serve as an alternative evaluation. Although it would be prohibitively expensive to measure $\mathcal{NC}$ on the vast and sometimes obscure pre-training data of most frontier production LLMs, doing so on a small set of unseen data (i.e. test data) would be realistic.
@@
At a high level, the number and density of clusters for a token can reflect its learned meanings and uses. This would be particularly useful as LLMs adapt to ever-evolving language and further expansion into non-English domains. Our formulae (Section 3) and results (Section 4) expose the pairwise token class interactions in noise, interference, classifier duality, and classifier agreement in the top-level features. Similarly to works in other domains [21, 55, 72, 115], these $\mathcal{NC}$ metrics can serve as a form of low-level interpretability to aid understanding certain behaviors of (L)LMs. Between tokens, one can often discern how related or interchangeable words are based on their pair-wise interactions, or how antithetical or unrelated they are based on orthogonality. For example, we present a rudimentary analysis of homonyms and English first names in Appendix Q.
@@
Foundation models are ubiquitous for their comprehensive capabilities and adaptability. As previous work discussed class imbalance [35, 58, 61, 62], our work may extend these strategies to measure and perhaps promote fairness in foundation LLMs, some of which are designed to be multilingual or multicultural. For example, [116] contemporarily explores the use of $\mathcal{UNC}3$ to mitigate biases in BERT-based [7] models.
While $\mathcal{NC}$ itself would not lead to unfairness, its potential interpretability may, in theory, enable an (adversarial) agent to measure and optimize for (un)fairness as they manipulate an LLM.
@@ 8 Conclusion
In this paper, we apply the neural collapse ($\mathcal{NC}$) framework to the next-token prediction problem, where models are undertrained and next-tokens are variably drawn from numerous and imbalanced token classes. We leverage canonical and more recent metrics to demonstrate that $\mathcal{NC}$ emerges as we scale the size and training of hundreds of causal language models. Our results show a correlation between $\mathcal{NC}$ and generalization, a relationship that persists even when the model scale is fixed.
In the short term, our work presents rudimentary techniques to analyze and interpret token-level properties of (L)LMs. We anticipate future work to suitably adapt $\mathcal{NC}$ (and related frameworks) to the still-fresh frontier of autoregressive language modeling. Researchers could then effectively capitalize on previous learnings from $\mathcal{NC}$ to better understand and improve the pre/post-training processes of increasingly complex and large language (multimodal) models.
Acknowledgements
We thank Elliot Creager, David Glukhov, Daniel Johnson, Jivan Waber, and Colin Raffel for their helpful feedback and stimulating discussions. Aditya Mehrotra and Yu Bo Gao provided technical assistance in our implementations. We acknowledge the support of Canada’s Natural Sciences and Engineering Research Council (NSERC, www.nserc-crsng.gc.ca/). This research was enabled in part by resources from Calcul Québec (www.calculquebec.ca), the Digital Research Alliance of Canada (www.alliancecan.ca), and the Vector Institute (www.vectorinstitute.ai).
Appendix A Dataset
TinyStories is a dataset of short children’s stories generated by GPT-3.5 and GPT-4 [2], released with the CDLA-Sharing-1.0 licence. We trained and evaluated on the first version, as described:
  • Their experimental setup [2] called for the GPT-2 [68] tokenizer, of which only a subset vocabulary $\mathbb{V}=\llbracket 1,29233\rrbracket$ appears in TinyStories.
[…]
Figure 4: The $500$ most frequent classes from TinyStories [2] exhibit significant sample imbalance. Despite the synthetic nature of TinyStories, such a distribution is typical of natural language [11, 12].
A.1 Alternative (Real) Datasets
The study of $\mathcal{NC}$ in causal language modeling at the token level would be very expensive, so the motivation to use a small dataset is clear. However, most commonly used text datasets such as WikiText [117], BookCorpus [118], CommonCrawl [119], or most subsets from the Pile [120] are much too complex and broad to be effectively compressed by CLMs of the scale that we work with.
WikiText-2 and WikiText-103 present significant drawbacks for our experiments. Both datasets contain a considerable amount of low-quality data that does not concentrate on essential linguistic structures such as grammar, vocabulary, facts, and reasoning. WikiText-2 has a similar empirical vocabulary to TinyStories under the GPT-Neo [80] tokenizer (27K vs. 29K) but only has around 34K rows of training data compared to 2.1M in TinyStories. Our small-scale NC experiment on WikiText-2 revealed that the models were very brittle and prone to overfitting. On the other hand, WikiText-103 is comparably sized to TinyStories but utilizes around 44K unique tokens. Our CLMs trained on WikiText-103 struggled to produce coherent sentences, likely due to the excessive breadth and information, as noted by the authors of TinyStories. Beyond these two, we were unable to find any real datasets that both followed established scaling laws [14, 15] for CLMs at our scale and are simple enough to suit the analysis of $\mathcal{NC}$.
A.2 On the Use of TinyStories
According to its authors, TinyStories [2] is explicitly designed to preserve the essential elements of natural language, such as grammar, vocabulary, facts, and reasoning, while being smaller and more refined in terms of its breadth and diversity. Unlike large corpora that can overwhelm small language models (SLMs) due to their excessive breadth and diversity, TinyStories offers a concentrated dataset that hones in on core linguistic structures and reasoning capabilities. This is evident in its small word vocabulary, consisting of approximately 1500 words that a child would use, and in its 29K empirical vocabulary under the GPT-Neo tokenizer.
Despite its concentrated nature, TinyStories enables models trained on it to produce grammatically correct, factual, and reasonable stories. Additionally, these models can be finetuned on specific instructions found in the TinyStories-Instruct dataset. The authors of TinyStories also demonstrate that their models can creatively produce stories that are dissimilar enough to their training data, indicating a balanced capability for generalization and creativity.
One particular advantage of TinyStories is the small vocabulary relative to total training tokens, rendering a reasonable number of classes with higher average token counts. This is relevant because the possibility of $\mathcal{NC}$ and a CLM’s ability to compress language data into distinct geometries depend partially on the ratios between embedding dimension, vocabulary size, and average token frequency. Conveniently, frequency analysis of the overall dataset produced a distribution (Figure 4) similar to real human language, so TinyStories should provide a good balance for an initial study of this $\mathcal{NC}$.
Additionally, TinyStories has a more regular structure as GPT-3.5/4 was instructed to produce children’s stories with certain themes and forms with a conservative vocabulary. We believe this would reduce the amount of clustering noise from the breadth of information and structures in real general data, and allow our smaller CLMs to exhibit some clear trends toward $\mathcal{NC}$.
Furthermore, TinyStories was created using GPT 3.5/4, which are advanced language models with significantly larger architectures trained on orders of magnitude more tokens, helping minimize the effect of the synthetic nature of the generated dataset. We also considered a possible effect of model collapse as a result of training on synthetic data [121] and follow-up work [122] suggest that a single iteration of data generation (as generated TinyStories) has a very negligible model collapse.
Appendix B Model Architectural Details
@@ Appendix C Optimization
Figure 5: Average (logarithmic) class-distance normalized variance (CDNV, $\mathcal{NC}_{1}$) (left) and validation (cross-entropy) loss (right) with respect to training epochs.
Appendix D Embeddings Collection & $\mathcal{NC}$ Analysis
Codes for (post-)training analysis are hosted on GitHub:
One pass over the train set for embeddings collection can take up to 6 hours on a single NVIDIA A100 (40GB) GPU. Analysis of a single metric for a given model takes no longer than 5 minutes.
Appendix E Within-Class Variability Collapse with Scale – $\mathcal{NC}_{1}$
Figure 6: Average (logarithmic) class-distance normalized variance (CDNV) is reduced ($\mathcal{NC}_{1}$) when scaling width ($d$) and across training for 1 (left) through 10 (right) epochs with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Figure 7: Average (logarithmic) class-distance normalized variance (CDNV) is reduced ($\mathcal{NC}_{1}$) when scaling depth ($L$) and across training for 1 (left) through 10 (right) epochs with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Appendix F Mean Norms Growth with Scale – (Related to $\mathcal{NC}_{2}$)
@@
Appendix G Equinormness with Scale – $\mathcal{(G)NC}_{2}$
Figure 10: Variation in embedding norms decreases ($\mathcal{NC}_{2}$) when scaling width ($d$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom). Note that the degree of equinormness eventually plateaus for sufficiently deep and trained models.
Figure 11: Variation in embedding norms decreases ($\mathcal{NC}_{2}$) when scaling depth ($L$) in models trained for 1 (left) through 10 (right) with weight decays $\beta=0.0005$ (top) and $\beta=0.1$ (bottom).
Appendix H Interference with Scale – (Related to $\mathcal{NC}_{2}$)